### v2

@@ 3.1 Comparing time streams
Figure 4 compares mock time streams with MUSTANG-2 observations, revealing a close resemblance between the modeled and actual time streams. The left panel shows the power spectra of the time streams depicted on the right. Prior to computing the power spectra, the time series are apodized with a tapered cosine window.
The power spectra reveal three different regimes where the atmosphere (left), pink (middle), and white noise (right) are dominant as a function of the sky sampling frequency. We found that a white noise level of TK (corrected for the calibration parameter and the opacity of the atmosphere, see Eq. 5) was required to reproduce the real power spectrum. The -term was also added in order to reproduce the real-time streams.
@@ 3.2 Comparing noise maps
Transforming a synthetic time stream into a map of the sky is an art of its own. As maria is designed specifically for generating synthetic time streams, we also implemented a simple, easy-to-use Python-based mapmaking tool into the routine, which we based on MIDAS, the MUSTANG IDL Data Analysis System666MIDAS when fully unraveled stands for the MUltiplexed Superconducting Quantum Interface Device Transition Edge Sensor Array at Ninety Gigahertz Interactive Data Language Data Analysis System. (see Romero et al., 2020, and references therein). But as maria is a generic tool that can simulate various different telescopes and observation strategies, one might not always want to use the same mapmaker. Hence, we also made the TOD exportable to FITS tables and Hierarchical Data Format (HDF5) files so that one can use whichever mapmaker they prefer.
@@
Even though the fake and real scans show similar characteristics and sensitivities, we want to point out several second-order effects that can still cause minor differences between the real and mock data. The most notable one is that on timescales of an hour, the weather at the Green Bank site might vary, for instance, becoming cloudy. This can worsen or suddenly improve weather conditions, which might not be modeled in maria and lead to variations in the effective beam size. Additionally, surface deformations of MUSTANG-2’s primary dish change the effective shape of the beam, which needs to be accurately calibrated. Further, the scanning strategy maria adopts is not exactly the same as that used by the GBT in MUSTANG-2 observations, as the real one features a less regular scanning trajectory.
Further, our model for the noise components is likely an oversimplification, as correlations in the instrumental noise are not fully accounted for in the maria simulations. A better characterization of which part of the optical/signal path (e.g., amplifiers, readouts, detector thermal drifts, optical loading due to spillover, or surface deformations) causes which noise contribution in the noise power spectra could help improve the noise model when making maps. Finally, the real observations include contamination in the form of spikes at high sampling frequencies () that could originate from the pulse tube or other vibrating objects mounted on the GBT such as the feedarm.
@@ 3.3 Forecasting MUSTANG-2 observations
To further quantify the scale dependence of how information is lost due to filtering and processing, we compute the transfer function of the filtered map with respect to the input map (see Figure 9). Here, the transfer function is defined by taking the square root of the power spectra of both a noiseless and noisy 45-minute-long mock observation and dividing it by the power spectra of the input map. To prevent the power spectra from blowing up at large wave numbers (smaller scales), we added white noise to the image at a level of a tenth of the measured sensitivity in the central region in Figure 8. The power spectra are estimated by binning the Fourier transform of the map in log-spaced bins from and taking the average of the amplitude of the Fourier transform per -bin. The “noiseless” curve shows the ideal transfer function for the observations. It exhibits a flat -dependence up to , after which it drops to by , which is similar to what is derived from the fidelity map shown in Figure 7 and agrees well with real MUSTANG-2 observations (see e.g., Romero et al., 2020)
@@ 4.1 Simulation set up
AtLAST, as mentioned in Section 2.1.2, is an ongoing design study for a 50-meter single-dish facility to be located in the Atacama desert near ALMA. AtLAST’s concept is that of a facility observatory hosting up to six massive instruments and serving a wide range of science cases. For the purpose of this work, we focus on a particular observing mode, class of instrumentation, and science goal, but we note that future works will explore AtLAST’s observing capabilities more broadly as well as make forecasts for its scientific applicability. We note here that the telescope will have a 4.6-meter focal plane, corresponding to an instantaneous field of view of 2, and will allow extremely fast scan velocities of up to (e.g., in the middle of an azimuthal scan), and an acceleration of up to as defined in Klaassen et al. (2020).
@@
To Nyquist sample the beam, which has a FWHM for a 50-meter telescope at 93 GHz and assuming an average scanning velocity of /s, we need to scan with a read-out rate 90 that of MUSTANG-2, resulting in a sampling of 900 Hz. AtLAST will also observe at high frequencies up to GHz, increasing the rate to kHz. Furthermore, to maintain the same sky sampling, the number of detectors increases with the frequency squared if the FoV size remains constant. For the 93 GHz band, we need detectors to fill the FoV with the detectors placed at 1. Hence, for the high-frequency (GHz), AtLAST would need 20 million detectors to fill the FoV. Fortunately, as noted in Klaassen et al. (2020), the number of detectors increases by roughly a factor of ten every seven years, implying that instruments comprising of order 1 million detectors could be expected by 2030. However, both the high sampling rate and the number of detectors in the array result in stupendously large data rates. Simulating this with maria is thus computationally expensive and requires a large working memory (TB). The MUSTANG-2 mock observations can be run on a 2019 MacBook Pro with a 2.6 GHz 6-Core Intel Core i7 processor, but with the current implementation of maria (v1.0), an even bigger server (e.g., a 64-core workstation with 1 TB of random access memory) is not enough to mimic AtLAST-like configurations.
@@
To temporarily overcome the data rate problem, we simulate here an instrument with a reduced FoV of only () and limit the maximum scanning velocity to , which occurs at the center of the daisy scan. This will limit the number of detectors to 3000 (a factor of 32 smaller than AtLAST could host) and lower the data rate by sampling at 225 Hz (12 times lower). These adjustments make the data rates feasible to run with maria as we continue to improve the efficiency of the code. For the remainder of this work, the observations are set to the following configuration: it will point at a source in the position of an RA and Dec of at the date of 2022-08-01T23:00:00 which results in an average elevation of 57.3. The default scanning radius of the daisy scan will be the size of the FoV. Furthermore, we will simulate one band for all detectors centered on 92 GHz with a bandwidth of 52 GHz, which is ideal for detecting the SZ effect at those frequencies. Finally, we will simulate the atmosphere to a scale height of 3000 meters above Chajnantor and use an RMS fractional variation in PWV of 5% to get a total temperature of around which is a typical atmospheric value (Pardo et al., 2001) when observing at the APEX, ALMA, or future AtLAST site.
@@
Further, we adopt the measured mean sensitivity of the PA6 array ( at GHz, Naess et al. in prep.) of ACT as a white noise level for each detector used in our setup. As the ACT detector noise performance is nearly background-limited, better sensitivities per detector per time stamp cannot easily be achieved. Further, we normalized the ACT sensitivity with the ratio of the different bandpasses and corrected for the optical efficiency we simulated; Here, we assume similar optical efficiencies as for the GBT and MUSTANG-2. The resulting white noise level is four times lower than that of MUSTANG-2. Then, we scaled the pink noise level down from what we used in the MUSTANG-2 case with the same amplitude as the difference in the white noise levels. All this results in an antenna temperature of around after correcting the time streams for the optical efficiency and the atmosphere’s opacity. The resulting power spectrum and noise map are shown in Figure 10.
Before making this map, we highpass filter the time streams in Fourier space to remove scales larger than the field of view (i.e. Hz in the time domain, depending on the scanning speed). We then remove the first principal component, as is often done with real observations. The Fourier filtering removes the peaks caused by the scanning harmonics seen below 0.5 Hz in the upper panel of Figure 10. The Fourier filtering also removes most of the atmospheric contributions from the time streams.
@@ 4.2.1 Large spatial scales
In addition to collective area and resolution, one of the advantages of constructing a large (meter) single-dish telescope, such as AtLAST or CMB-HD (Sehgal et al., 2019; The CMB-HD Collaboration et al., 2022), turns out to be the enhanced correlation of atmospheric effects across beams of adjacent detectors, which facilitates better removal of the atmospheric signal.
@@
The Rayleigh height sets the boundary between the near and the far-field. From Eq. (7), we infer that in physical units (m), the near field beam behaves as a cylindrical column that transverses through the sky and diverges only at higher altitudes for larger dish sizes, as illustrated in Figure 11. Based on the geometry of the beam, we infer that a large dish sees more of the same atmosphere between two beams than a 6-m or a 42-cm aperture, which are the apertures sizes of the large aperture telescope (LAT) and small aperture telescope (SAT) used in Simons Observatory (SO, Ade et al., 2019). To illustrate, at a scale height of 1 km and adjacent detectors at 150 GHz spaced apart, with AtLAST, one detector’s beam would be shifted only by 48 mm relative to the next adjacent beam (i.e. roughly one-thousandth of the 50 meter column of atmosphere probed in the near field). This implies that one detector sees 99.1% of the same of the atmosphere as its neighbor. For the 6-m apertures, the same considerations imply the shift would be 87.9%, making time-lagged atmospheric signal removal more difficult. For the 42-cm, this would be even worse; however, as Figure 11 shows, a 42-cm aperture results in an extremely short near field, with the beams beginning to diverge around 69 meters, thus increasing the overlap in the instantaneous volumes of the atmosphere probed, which also should facilitate better atmospheric removal in post-processing.
@@
A fundamental assumption here is that a higher level of common mode signal should allow for better mitigation of the atmosphere, leading to the recovery of lower . This simulation shows that a 50 m dish could better recover signals at lower than 6 m or 42-cm sized apertures for a similar FoV. ACT has already demonstrated with a 6 meter aperture and 1 FoV the ability to recover scales down to (see, e.g., Coulton et al., 2023; Qu et al., 2023). Therefore, even if one assumes that the spatial scales one can recover are limited to scales less than or equal to the size of the FoV, this implies that AtLAST could also carry out CMB observations on similar or larger angular scales than the current CMB experiments.
@@ 4.2.2 Small spatial scales
Regarding the sensitivity at beam scales (high ), we can compare maria simulations with the predictions from the AtLAST sensitivity calculator.777See https://senscalc.atlast.uio.no/ There, the point source sensitivity is estimated by using the CASA corruption model set to Total Power mode (CASA Team et al., 2022) and scaling up the diameter of a single ALMA antenna to 50 meters. The sensitivity calculator is appropriate for maps of the line emission using state-of-the-art heterodyne instruments, which trade off a worse noise performance with respect to a bolometer to preserve phase coherence and achieve much higher spectral resolution (). However, this model does not account for correlated atmosphere, nor are scanning patterns implemented. Hence, using the sensitivity calculator to make forecasts for bolometric (e.g., broadband continuum) instruments is not ideal as it will act as if it has a single detector on the sky that will stare at the celestial object. Regardless, comparing the AtLAST sensitivity calculator with what maria outputs is a useful cross-check on results from maria.
The lower panel in Figure 10 displays the noise map for our simulated AtLAST observation, but with additional Fourier filtering and a common mode subtraction applied. Within the central region, this map has a beam sensitivity of 12.0 Jy/Beam when assuming a diffraction-limited beam shape. In contrast to the AtLAST sensitivity estimator, which provides a sensitivity of Jy/Beam for an 8.6 minute long observation scan888We used the average elevation of , dual polarization mode, and a HO profile percentile of 25., we find a slightly better per beam sensitivity (i.e., lower RMS) than theoretical predictions even without using the full AtLAST design specifics, and even though the 8.6-minute long integration time is distributed over the map, since the detector array moves on and off source as, for instance, it was for MUSTANG-2 (see Fig 2). Of course, the overall performance of maria depends somewhat on the actual instrument performance, and here we have conservatively adopted the same white and pink noise levels demonstrated by ACT. Regardless, the similar point source sensitivities appear to indicate that the two methodologies broadly agree.

### v3

@@ 3.1 Comparing time streams
Figure 4 compares mock time streams with MUSTANG-2 observations, revealing a close resemblance between the modeled and actual time streams. The time streams have a duration of minutes, which is the nominal duration of a single GBT/MUSTANG-2 scan. The left panel shows the power spectra of such a scan, which is also depicted on the right. Prior to computing the power spectra, the time series are apodized using a tapered cosine window. The power spectra reveal three different regimes where the atmosphere (left), pink (middle), and white noise (right) are dominant as a function of the sky sampling frequency. We found that a white noise level of TK (corrected for the calibration parameter and the opacity of the atmosphere, see Eq. 5) was required to reproduce the real power spectrum. The -term was also added in order to reproduce the real-time streams.
@@
Figure 6 shows the mean power spectra of all detectors from 13 consecutive real 8.6-minute long MUSTANG-2 scans, totaling approximately 2 hours of observation. The alignment of all power spectra indicates consistent statistical properties across scans. Therefore, the comparison between the simulated and real observations presented in Figure 4 also applies to longer integration times.
@@ 3.2 Comparing noise maps
3.2.1 Naive Mapmaking
Transforming a synthetic time stream into a map of the sky is an art of its own. As maria is designed specifically for generating synthetic time streams, we also implemented a simple, easy-to-use Python-based mapmaking tool into the routine, which we based on MIDAS, the MUSTANG IDL Data Analysis System777MIDAS when fully unraveled stands for the MUltiplexed Superconducting Quantum Interface Device Transition Edge Sensor Array at Ninety Gigahertz Interactive Data Language Data Analysis System. (see Romero et al., 2020, and references therein). But as maria is a generic tool that can simulate various different telescopes and observation strategies, one might not always want to use the same mapmaker. Hence, we also made the TOD exportable to FITS tables and Hierarchical Data Format (HDF5) files so that one can use whichever mapmaker they prefer.
@@ 3.2.1 Naive Mapmaking
3.2.2 Limitations of the simulated observations
Despite the similarity displayed by Figure 7, in this section, we discuss several second-order effects that can still cause minor differences between the real and mock data. The most notable one is that on timescales of several hours, the weather at the Green Bank site might vary, for instance, becoming cloudy. This can worsen or suddenly improve weather conditions, which might not be modeled in maria and lead to variations in the effective beam size. Additionally, surface deformations of MUSTANG-2’s primary dish change the effective shape of the beam, which needs to be accurately calibrated, as well as inaccuracies in the bandpass, which are not modeled in full detail in maria. Further, the scanning strategy maria adopts is not exactly the same as that used by the GBT in MUSTANG-2 observations, as the real one features a less regular scanning trajectory. The real observations also include contamination in the form of spikes at high sampling frequencies () that could originate from the pulse tube or other vibrating objects mounted on the GBT, such as the feed arm.
Noteworthy as well; our model for the noise components is likely an oversimplification, as correlations in the instrumental noise are not fully accounted for in the maria simulations. Further, the MUSTANG-2 detectors are not background-limited, meaning that we had to manually adjust the level of white noise in the simulations to match the observations in Figure 4. Although maria has the option to use background-limited estimates of the white noise based on the implementation of Bryan et al. (2018), a better characterization of which part of the optical/signal path (e.g., amplifiers, readouts, detector thermal drifts, optical loading due to spillover, or surface deformations) causes which noise contribution in the noise power spectra could help improve the noise model when making maps.
Finally, although the atmospheric component in maria is state-of-the-art, it has its limitations in accurately creating atmospheric corruption. maria adopts average quantities for typical minute-to-minute variations in PWV and the absolute scale height, which might vary over time (see, e.g., Morris et al., 2024, for a more detailed discussion). Further, there are unknowns, such as how the size of the boundary layer affects inhomogeneities in the atmosphere at different heights, how the local terrain influences local turbulence, and the effect of a non-constant horizontal wind velocity. However, as stated at the beginning of this section, these adjustments only have minor impact on the overall simulated atmosphere, as we show in Figures 4, 5, and 7, we can accurately retrieve the statistical behavior of the observations and obtain quantitatively similar noise maps.
@@ 3.3 Forecasting MUSTANG-2 observations
To further quantify the scale dependence of how information is lost due to filtering and processing, we compute the transfer function of the filtered map with respect to the input map (see Figure 10). Here, the transfer function is defined by taking the square root of the power spectra of the mock observation and dividing it by the power spectra of the input map (which is corrected for beam smearing). To prevent the power spectra from blowing up at large wave numbers (smaller scales), we added white noise to the image at a level of a tenth of the measured sensitivity in the central region in Figure 9. The power spectra are estimated by binning the Fourier transform of the map in log-spaced bins from and taking the average of the amplitude of the Fourier transform per -bin. The “noiseless” curve shows the ideal transfer function for the observations. It exhibits a flat -dependence up to , after which it drops to by , which is similar to what is derived from the fidelity map shown in Figure 8 and agrees well with real MUSTANG-2 observations (see e.g., Romero et al., 2020).
@@ 4.1 Simulation set up
AtLAST is a proposed 50-meter diameter single-dish facility to be built in the Atacama desert near ALMA. AtLAST has already undergone a 4-year EU-funded design study and has been selected for a new EU infrastructure development grant starting in 2025. AtLAST’s concept is that of a facility observatory hosting up to six massive instruments and serving a wide range of science cases. For the purpose of this work, we focus on a particular observing mode, class of instrumentation, and science goal, but we note that future works will explore AtLAST’s observing capabilities more broadly as well as make forecasts for its scientific applicability. We note here that the telescope will have a 4.7-meter focal plane, corresponding to an instantaneous field of view of 2∘, and will allow extremely fast scan velocities of up to (e.g., in the middle of an azimuthal scan), and an acceleration of up to as defined in Klaassen et al. (2020).
@@
To Nyquist sample the beam, which has a FWHM for a 50-meter telescope at 93 GHz and assuming an average scanning velocity of /s, we need to scan with a read-out rate 90 that of MUSTANG-2, resulting in a sampling of 900 Hz. AtLAST will also observe at high frequencies up to GHz, increasing the rate to kHz. Furthermore, to maintain the same sky sampling, the number of detectors increases with the frequency squared if the FoV size remains constant. For the 93 GHz band, we need detectors to fill the FoV with the detectors placed at 1. Hence, for the high-frequency (GHz), AtLAST would need 20 million detectors to fill the FoV. Fortunately, as noted in Klaassen et al. (2020), the number of detectors increases by roughly a factor of ten every seven years, implying that instruments comprising of order 1 million detectors could be expected by 2030. However, both the high sampling rate and the number of detectors in the array result in stupendously large data rates. Simulating this with maria is thus computationally expensive and requires a large working memory (TB). The MUSTANG-2 mock observations finished its run within five minutes on a 2019 MacBook Pro with a 2.6 GHz 6-Core Intel Core i7 processor, but with the current implementation of maria (v1.0), an even bigger server (e.g., a 64-core workstation with 1 TB of random access memory) is not enough to mimic AtLAST-like configurations.
@@
To temporarily overcome the data rate problem, we simulate here an instrument with a reduced FoV of only () and limit the maximum scanning velocity to , which occurs at the center of the daisy scan. This will limit the number of detectors to 3000 (a factor of 32 smaller than AtLAST could host) and lower the data rate by sampling at 225 Hz (12 times lower). These adjustments make the data rates feasible to run with maria as we continue to improve the efficiency of the code. For the remainder of this work, the observations are set to the following configuration: it will point at a source in the position of an RA and Dec of at the date of 2022-08-01T23:00:00 which results in an average elevation of 57.3∘. The default scanning radius of the daisy scan will be the size of the FoV. Furthermore, we will simulate one band for all detectors centered on 92 GHz with a bandwidth of 52 GHz, which is ideal for detecting the SZ effect at those frequencies. Finally, we will simulate the atmosphere to a scale height of 3000 meters above Chajnantor and use an RMS fractional variation in PWV of 5% to get a total temperature of around which is a typical atmospheric value (Pardo et al., 2001) when observing at the APEX, ALMA, or future AtLAST site.
@@
Further, we adopt the measured mean sensitivity of the PA6 array ( at GHz, Naess et al. in prep.) of ACT as a white noise level for each detector used in our setup. As the ACT detector noise performance is nearly background-limited, better sensitivities per detector per time stamp cannot easily be achieved. Further, we normalized the ACT sensitivity with the ratio of the different bandpasses and corrected for the optical efficiency we simulated; Here, we assume similar optical efficiencies as for the GBT and MUSTANG-2. The resulting white noise level is four times lower than that of MUSTANG-2. Then, we scaled the pink noise level down from what we used in the MUSTANG-2 case with the same amplitude as the difference in the white noise levels. All this results in an antenna temperature of around after correcting the time streams for the optical efficiency and the atmosphere’s opacity. The resulting power spectrum and noise map of an 8.6-minute long scan are shown in Figure 11.888Adopting an 8.6-minute scan duration for AtLAST is arbitrary and was chosen solely for consistency with the MUSTANG-2 comparison. The scan duration for AtLAST is not fixed and will likely depend on the specific instrument and scientific goals. Before making this map, we highpass filter the time streams in Fourier space to remove scales larger than the field of view (i.e., Hz in the time domain, depending on the scanning speed). We then remove the first principal component, as is often done with real observations. The Fourier filtering removes the peaks caused by the scanning harmonics seen below 0.5 Hz in the upper panel of Figure 11. The Fourier filtering also removes most of the atmospheric contributions from the time streams.
@@ 4.2.1 Large spatial scales
In addition to collecting area and resolution, one of the advantages of constructing a large (meter) single-dish telescope, such as AtLAST or CMB-HD (Sehgal et al., 2019; The CMB-HD Collaboration et al., 2022), turns out to be the enhanced correlation of atmospheric effects across beams of adjacent detectors, which facilitates better removal of the atmospheric signal.
@@
The Rayleigh height sets the boundary between the near and the far-field. From Eq. (7), we infer that in physical units (m), the near field beam behaves as a cylindrical column that transverses through the sky and diverges only at higher altitudes for larger dish sizes, as illustrated in Figure 12. Based on the geometry of the beam, we infer that a large dish sees more of the same atmosphere between two beams than a 6-m or a 42-cm aperture, which are the apertures sizes of the large aperture telescope (LAT) and small aperture telescope (SAT) used in SO. To illustrate, at a scale height of 1 km and adjacent detectors at 150 GHz spaced apart, with AtLAST, one detector’s beam would be shifted only by 48 mm relative to the next adjacent beam (i.e. roughly one-thousandth of the 50 meter column of atmosphere probed in the near field). This implies that one detector sees 99.1% of the same of the atmosphere as its neighbor. For the 6-m apertures, the same considerations imply the shift would be 87.9%, making time-lagged atmospheric signal removal more difficult. For the 42-cm, this would be even worse; however, as Figure 12 shows, a 42-cm aperture results in an extremely short near field, with the beams beginning to diverge around 69 meters, thus increasing the overlap in the instantaneous volumes of the atmosphere probed, which also should facilitate better atmospheric removal in post-processing.
@@
A fundamental assumption here is that a higher level of common mode signal should allow for better mitigation of the atmosphere, leading to the recovery of lower . This simulation shows that a 50 m dish could better recover signals at lower than 6 m or 42-cm sized apertures for a similar FoV. ACT has already demonstrated with a 6 meter aperture and 1∘ FoV the ability to recover scales down to (see, e.g., Coulton et al., 2024; Qu et al., 2024) and SPT probing down to (Henning et al., 2018). Therefore, even if one assumes that the spatial scales one can recover are limited to scales less than or equal to the size of the FoV, this implies that AtLAST could also carry out CMB observations on similar or larger angular scales than the current CMB experiments.
@@ 4.2.2 Small spatial scales
Regarding the sensitivity at beam scales (high ), we can compare maria simulations with the predictions from the AtLAST sensitivity calculator.999See https://senscalc.atlast.uio.no/ There, the point source sensitivity is estimated by using the CASA corruption model set to Total Power mode (CASA Team et al., 2022) and scaling up the diameter of a single ALMA antenna to 50 meters. The sensitivity calculator is appropriate for maps of the line emission using state-of-the-art heterodyne instruments, which trade off a worse noise performance with respect to a bolometer to preserve phase coherence and achieve much higher spectral resolution (). However, this model does not account for correlated atmosphere, nor are scanning patterns implemented. Hence, using the sensitivity calculator to make forecasts for bolometric (e.g., broadband continuum) instruments is not ideal as it will act as if it has a single detector on the sky that will stare at the celestial object. Regardless, comparing the AtLAST sensitivity calculator with what maria outputs is a useful cross-check on results from maria.
The lower panel in Figure 11 displays the noise map for our simulated AtLAST observation, but with additional Fourier filtering and a common mode subtraction applied. Within the central region, this map has a beam sensitivity of 12.0 Jy/Beam when assuming a diffraction-limited beam shape. In contrast to the AtLAST sensitivity estimator, which provides a sensitivity of Jy/Beam for an 8.6 minute long observation scan101010We used the average elevation of , dual polarization mode, and a H2O profile percentile of 25., we find a slightly better per beam sensitivity (i.e., lower RMS) than theoretical predictions even without using the full AtLAST design specifics, and even though the 8.6-minute long integration time is distributed over the map, since the detector array moves on and off source as, for instance, it was for MUSTANG-2 (see Fig 2). Of course, the overall performance of maria depends somewhat on the actual instrument performance, and here we have conservatively adopted the same white and pink noise levels demonstrated by ACT. Regardless, the similar point source sensitivities appear to indicate that the two methodologies broadly agree.